# Smoothing Variances Across Time: Adaptive Stochastic Volatility

### v4

@@ A.2 Theorem 2
$\prod_{h=0}^{\infty}\cos(\frac{\pi t}{2^{h}})$ $=\prod_{h=0}^{\infty}\frac{\sin(\frac{\pi t}{2^{h-1}})}{2\sin(\frac{\pi t}{2^{h}})}=\lim_{n\to\infty}\prod_{h=0}^{n}\frac{\sin(\frac{\pi t}{2^{h-1}})}{2\sin(\frac{\pi t}{2^{h}})}$
@@
$\prod_{h=0}^{\infty}\sec(\frac{\pi t}{2^{h}})=\frac{1}{\prod_{h=0}^{\infty}\cos(\frac{\pi t}{2^{h}})}=\frac{2\pi t}{\sin(2\pi t)}=\Gamma(1-2t)\Gamma(1+2t)=\mathrm{B}(1-2t,1+2t)$
$\frac{2\pi t}{\sin(2\pi t)}=\Gamma(1-2t)\Gamma(1+2t)$ is by the reflection relation. $Logistic(\mu,s)$ has the the following moment generating function $\text{exp}(-\mu t)\mathrm{B}(1-st,1+st)$. Since the moment generating function uniquely determines the random variable, and by Theorem 2,
$z_{t}=\sum_{h=0}^{\infty}z_{t,h}\stackrel{{\scriptstyle a.s}}{{\rightarrow}}Logistic(0,2)$
@@ A.3 Theorem 3
Define $\lambda_{t}:=\text{exp}(v_{t})$. We showed that $f(\lambda_{t})=\frac{1}{(1+\lambda_{t})^{2}}$. We have:
[…]
@@
[…]
$=\frac{1}{2\sqrt{2\pi}}\text{exp}(\Delta^{2}h_{t}/2)E_{1}(\Delta^{2}h_{t}/2),$
where $E_{1}()$ is the exponential integral function, which satisfies the following upper and lower bound $\forall t>0$: exp(-t)2 log(1 + 2t) ¡ E_1(t) ¡ exp(-t) log(1 + 1t) Thus, f(Δh_t) ¡ 122 π log(1 + 2Δ2ht) Similarly for the lower bound, f(Δh_t) ¿ 182 π log(1 + 4Δ2ht) The lower and upper bound for $\Delta h_{t}\neq 0$ is shown. The lower bound clearly approaches infinity as $\Delta h_{t}$ approaches 0, which completes the proof.
@@ A.4 Theorem 4
To prove identifiability of the hierarchical model, we need to establish that
@@
By model specification $\boldsymbol{\beta}|\boldsymbol{\lambda}_{\beta}\sim N(\boldsymbol{0},\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T})$ and $\boldsymbol{h}|\boldsymbol{\lambda}_{h}\sim N(\boldsymbol{0},\mathbf{S}\Lambda_{h}\mathbf{S}^{T})$ where $\Lambda_{\beta}=diag(\boldsymbol{\lambda}^{2}_{\beta})$, $\Lambda_{h}=diag(\boldsymbol{\lambda}^{2}_{h})$ and $\mathbf{S}$ is the un-differencing operator matrix corresponding to the differencing order k in $\Delta^{k}\beta$ and $\Delta^{k}h$.
@@
$=\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg{\}}+\mathbf{S}\Lambda_{\beta}^{2}\mathbf{S}^{T}.$
@@
$\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg{\}}+\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T}=\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}^{*}\mathbf{S}^{T})\bigg{\}}+\mathbf{S}\Lambda_{\beta}^{*}\mathbf{S}^{T},$
$\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg{\}}-\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}^{*}\mathbf{S}^{T})\bigg{\}}=\mathbf{S}\Lambda_{\beta}^{*}\mathbf{S}^{T}-\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T}.$
@@ B.1 $j$
[…]
By Omori et al. [2007],
@@
which is what we wanted to show. Exact distribution on $j_{t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}Categorical(\pi^{\text{Omori}})$ as well as corresponding mean and the variance parameter of each component is described in Omori et al. [2007].
@@ B.3 $v$
We use the all-without-loop (AWOL) sampler by Kastner and Frühwirth-Schnatter [2014] with parameter expansion on the error term with scale mixture normal distribution for sampling $\textbf{v},\mu$ and $\phi.$ By the conditional independence, the likelihood with respect to $\boldsymbol{v}$, reduces to $f(\boldsymbol{\omega^{*}}|v)$ where $\omega_{1}^{*}=\text{log}(h_{1}^{2})$ and $\omega_{t}^{*}=\text{log}((h_{t}-h_{t-1})^{2}),\forall t\geq 2.$ With the parameter expansion on the likelihood with 10-component Gaussian Mixture proposed by Omori et al. [2007],
@@
$f(v^{*}_{1}|\xi_{1},\mu)$ $=f(v_{1}|\xi_{0},\mu)=\mathcal{N}(v^{*}_{t}|\mu,\frac{1}{\xi_{1}})$
$f(v^{*}_{t}|\xi_{t-1},\mu,\phi)$ $=f(v_{t}-\phi v_{t-1}|\xi_{t-1},\mu,\phi)=\mathcal{N}(v^{*}_{t}|\mu(1-\phi),\frac{1}{\xi_{t-1}})$ $t\geq 2,$
@@ B.5 $\xi$
Based on Polson et al. [2013], we have
@@
where $\mathcal{PG}$ represents the density function for Polya-Gamma random variable.
@@ B.7 $\xi_{\mu}$
Based on the scale mixture representation of the Z-distribution Polson et al. [2013]:
@@ B.8 $\phi$
Unlike other parameters, $\phi$ is not conditionally conjugate to its likelihood. Thus, the slice sampling method Neal [2003] is used for the MCMC sampling. By taking the linear transformation of $v_{t}$, Let’s consider its likelihood becomes $\hat{v}_{\mu}:=\frac{1}{T-1}\sum_{t=2}^{T}v_{\mu,t}$, where $v_{\mu,t}=\frac{1}{2}\bigg{(}\frac{v_{t}-\mu}{v_{t-1}-\mu}+1\bigg{)}$. We have the following conditional likelihood and prior for $\phi$:
$f(\hat{v}_{\mu}|\mu,\boldsymbol{\xi})=\mathcal{N}\bigg{(}\hat{v}_{\mu}\bigg{|}\frac{\phi+1}{2},\frac{1}{(T-1)^{2}}\sum_{t=2}^{T}\frac{1}{4\xi_{t-1}(v_{t-1}-\mu)^{2}}\bigg{)}$
$f(\frac{\phi+1}{2})\sim Beta\bigg{(}\frac{\phi+1}{2}\bigg{|}10,2\bigg{)}$

### v5

@@ A.2 Theorem 2
$\prod_{\ell=0}^{\infty}\cos(\frac{\pi t}{2^{\ell}})$ $=\prod_{\ell=0}^{\infty}\frac{\sin(\frac{\pi t}{2^{\ell-1}})}{2\sin(\frac{\pi t}{2^{\ell}})}=\lim_{n\to\infty}\prod_{\ell=0}^{n}\frac{\sin(\frac{\pi t}{2^{\ell-1}})}{2\sin(\frac{\pi t}{2^{\ell}})}$
@@
$\prod_{\ell=0}^{\infty}\sec(\frac{\pi t}{2^{\ell}})=\frac{1}{\prod_{\ell=0}^{\infty}\cos(\frac{\pi t}{2^{\ell}})}=\frac{2\pi t}{\sin(2\pi t)}=\Gamma(1-2t)\Gamma(1+2t)=\mathrm{B}(1-2t,1+2t)$
$(2\pi t)/\sin(2\pi t)=\Gamma(1-2t)\Gamma(1+2t)$ is by the reflection relation. $\text{Logistic}(\mu,s)$ has the the following moment generating function $\exp(-\mu t)\mathrm{B}(1-st,1+st)$. Since the moment generating function uniquely determines the random variable, and by Theorem 2,
$z_{t}=\sum_{\ell=0}^{\infty}z_{\ell,t}\stackrel{{\scriptstyle a.s}}{{\rightarrow}}\text{Logistic}(0,2)$
@@ A.3 Theorem 3
Define $\lambda_{t}:=\exp(v_{t})$. We showed that $f(\lambda_{t})=1/(1+\lambda_{t})^{2}$. We have:
$f(\Delta h_{t})=\int_{0}^{\infty}\frac{1}{\sqrt{2\pi\lambda_{t}^{2}}}\text{exp}\bigg(-\frac{(\Delta h_{t})^{2}}{2\lambda_{t}^{2}}\bigg)\frac{1}{(1+\lambda_{t})^{2}}d\lambda_{t}$
@@
$f(\Delta h_{t})$ $\leq\int_{0}^{\infty}\frac{1}{\sqrt{2\pi\lambda_{t}^{2}}}\text{exp}\bigg(-\frac{(\Delta h_{t})^{2}}{2\lambda_{t}^{2}}\bigg)\frac{1}{(1+\lambda_{t}^{2})}d\lambda_{t}$
$=\frac{1}{2\sqrt{2\pi}}\int_{0}^{\infty}\frac{1}{1+u}\text{exp}\bigg(-\frac{u(\Delta h_{t})^{2}}{2}\bigg)du$ $u=\frac{1}{\lambda_{t}^{2}}$
$=\frac{1}{2\sqrt{2\pi}}\exp\{(\Delta h_{t})^{2}/2\}E_{1}((\Delta h_{t})^{2}/2),$
where $E_{1}()$ is the exponential integral function, which satisfies the following upper and lower bound $\forall t>0$: exp(-t)2 log(1 + 2t) ¡ E_1(t) ¡ exp(-t) log(1 + 1t) Thus, f(Δh_t) ¡ 122 π log(1 + 2(Δht)2) Similarly for the lower bound, f(Δh_t) ¿ 182 π log(1 + 4(Δht)2) The lower and upper bound for $\Delta h_{t}\neq 0$ is shown. The lower bound clearly approaches infinity as $\Delta h_{t}$ approaches 0, which completes the proof.
@@ A.4 Theorem 4
Let $\boldsymbol{\beta}:=\{\beta_{1},\ldots,\beta_{T}\}$ and $\boldsymbol{h}:=\{h_{1},\ldots,h_{T}\}$ denote the latent mean and log-volatility processes, respectively. For notational convenience, define $\boldsymbol{\lambda}_{\beta}:=\{\sigma_{\beta,1}^{2},\ldots,\sigma_{\beta,T}^{2}\}$ and $\boldsymbol{\lambda}_{h}:=\{\sigma_{h,1}^{2},\ldots,\sigma_{h,T}^{2}\}$. To prove identifiability of the hierarchical model, we need to establish that
@@
By model specification $\boldsymbol{\beta}|\boldsymbol{\lambda}_{\beta}\sim N(\boldsymbol{0},\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T})$ and $\boldsymbol{h}|\boldsymbol{\lambda}_{h}\sim N(\boldsymbol{0},\mathbf{S}\Lambda_{h}\mathbf{S}^{T})$ where $\Lambda_{\beta}=diag(\boldsymbol{\lambda}^{2}_{\beta})$, $\Lambda_{h}=\text{diag}(\boldsymbol{\lambda}^{2}_{h})$ and $\mathbf{S}$ is the un-differencing operator matrix corresponding to the differencing order k in $\Delta^{k}\beta$ and $\Delta^{k}h$.
@@
$=\text{exp}\bigg\{\frac{1}{2}\text{diag}(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg\}+\mathbf{S}\Lambda_{\beta}^{2}\mathbf{S}^{T}.$
@@
$\text{exp}\bigg\{\frac{1}{2}\text{diag}(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg\}+\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T}=\text{exp}\bigg\{\frac{1}{2}\text{diag}(\mathbf{S}\Lambda_{h}^{*}\mathbf{S}^{T})\bigg\}+\mathbf{S}\Lambda_{\beta}^{*}\mathbf{S}^{T},$
$\text{exp}\bigg\{\frac{1}{2}\text{diag}(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg\}-\text{exp}\bigg\{\frac{1}{2}\text{diag}(\mathbf{S}\Lambda_{h}^{*}\mathbf{S}^{T})\bigg\}=\mathbf{S}\Lambda_{\beta}^{*}\mathbf{S}^{T}-\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T}.$
@@ B.1 $j$
$\mathrm{p}(j_{t}=k|h_{t},v_{t},s_{t},\xi_{t},\mu,\xi_{\mu},\phi,y_{t}^{*})$ $=\mathrm{p}(j_{t}=k|h_{t},y_{t}^{*})=\frac{f(y_{t}^{*}|h_{t},j_{t})\mathrm{p}(j_{t}=k)}{\int f(y_{t}^{*}|h,j_{t})f(j_{t})dj_{t}}$
By omorietal,
@@
which is what we wanted to show. Exact distribution on $j_{t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\text{Categorical}(\pi^{\text{Omori}})$ as well as corresponding mean and the variance parameter of each component is described in omorietal.
@@ B.3 $v$
We use the all-without-loop (AWOL) sampler by Kastner_2014 with parameter expansion on the error term with scale mixture normal distribution for sampling $\textbf{v},\mu$ and $\phi.$ By the conditional independence, the likelihood with respect to $\boldsymbol{v}$, reduces to $f(\boldsymbol{\omega^{*}}|v)$ where $\omega_{1}^{*}=\text{log}(h_{1}^{2})$ and $\omega_{t}^{*}=\text{log}((h_{t}-h_{t-1})^{2}),\forall t\geq 2.$ With the parameter expansion on the likelihood with 10-component Gaussian Mixture proposed by omorietal,
@@
$f(v^{*}_{1}|\xi_{1},\mu)$ $=f(v_{1}|\xi_{0},\mu)=\mathcal{N}(v^{*}_{t}|\mu,1/\xi_{1})$
$f(v^{*}_{t}|\xi_{t-1},\mu,\phi)$ $=f(v_{t}-\phi v_{t-1}|\xi_{t-1},\mu,\phi)=\mathcal{N}(v^{*}_{t}|\mu(1-\phi),1/\xi_{t-1})$ $t\geq 2,$
@@ B.5 $\xi$
Based on polyagamma, we have
@@
where $\mathcal{PG}$ represents the density function for Pólya-Gamma random variable.
@@ B.7 $\xi_{\mu}$
Based on the scale mixture representation of the Z-distribution polyagamma:
@@ B.8 $\phi$
Prior distribution is imposed on $\phi^{*}:=(\phi+1)/2$, where $\phi^{*}\sim\text{Beta}(1/2,1/2)$. $\phi=2\phi^{*}-1$. We may rewrite the equation as a regression form:
$v_{t}=\mu+\phi(v_{t-1}-\mu)+\eta_{t},$
$v_{t}=\mu+(2\phi^{*}-1)(v_{t-1}-\mu)+\eta_{t},$
$(v_{t}+v_{t-1}-2\mu)=\phi^{*}(2v_{t-1}-2\mu)+\eta_{t}.$
Because $\eta_{t}$ is conditionally normal, the formulation can be viewed as a simple linear regression where the dependent variable is $v_{t}+v_{t-1}-2\mu$, the regressor is $(2v_{t-1}-2\mu)$, and the regression coefficient $\phi^{*}$. Conditional on all other parameters, the likelihood for $\phi^{*}$ is therefore univariate Gaussian. Combining this likelihood with the non-conjugate, but exponential-family $\text{Beta}(1/2,1/2)$ prior yields a smooth posterior over $(0,1)$, which is sampled using slice sampling algorithm by slice.
Appendix C Simulation Set-Up
DGP1: SV with 1 Regime DGP2: SV with 2 Regimes
$y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=-7.89+0.94(h_{t-1}-7.89)+0.327u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=m_{s_{t}}+0.85(h_{t-1}-m_{s_{t-1}})+0.0461u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $m_{s_{t}}=\begin{cases}\,3.35\,,&\text{if $s_{t}=0$.}\\ \,10.47\,,&\text{if $s_{t}=1$.}\end{cases}$
DGP3: SV with 3 Regimes DGP4: GARCH with 1 Regime
$y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=m_{s_{t}}+0.65(h_{t-1}-m_{s_{t-1}})+0.065u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $m_{s_{t}}=\begin{cases}\,1.46\,,&\text{if $s_{t}=0$.}\\ \,3.91\,,&\text{if $s_{t}=1$.}\\ \,6.22\,,&\text{if $s_{t}=2$.}\end{cases}$ $y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=0.001+0.263y_{t-1}^{2}+0.705\sigma_{y,t-1}^{2}$
DGP5: GARCH with 2 Regimes DGP6: GARCH with 3 Regimes
$y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=m_{s_{t}}+\alpha_{s_{t}}y_{t-1}^{2}+\beta_{s_{t}}\sigma_{y,t-1}^{2}$ $m_{s_{t}}=\begin{cases}\,1.671\,,&\text{if $s_{t}=0$.}\\ \,108.67\,,&\text{if $s_{t}=1$.}\end{cases}$ $\alpha_{s_{t}}=\begin{cases}\,0.0003\,,&\text{if $s_{t}=0$.}\\ \,0.0181\,,&\text{if $s_{t}=1$.}\end{cases}$ $\beta_{s_{t}}=\begin{cases}\,0.99\,,&\text{if $s_{t}=0$.}\\ \,0.9042\,,&\text{if $s_{t}=1$.}\end{cases}$ $y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=m_{s_{t}}+\alpha_{s_{t}}y_{t-1}^{2}+\beta_{s_{t}}\sigma_{y,t-1}^{2}$ $m_{s_{t}}=\begin{cases}\,0.002\,,&\text{if $s_{t}=0$.}\\ \,0.051\,,&\text{if $s_{t}=1$.}\\ \,13.35\,,&\text{if $s_{t}=2$.}\end{cases}$ $\alpha_{s_{t}}=\begin{cases}\,0.073\,,&\text{if $s_{t}=0$.}\\ \,0.0035\,,&\text{if $s_{t}=1$.}\\ \,0.102\,,&\text{if $s_{t}=2$.}\end{cases}$ $\beta_{s_{t}}=\begin{cases}\,0.927\,,&\text{if $s_{t}=0$.}\\ \,0.906\,,&\text{if $s_{t}=1$.}\\ \,0.895\,,&\text{if $s_{t}=2$.}\end{cases}$
DGP7: Sinusoidal DGP8: Piecewise Constant
$y_{t}\sim\mathcal{N}(0,\exp(h_{t}))$ $h=A\sin(10(2\pi t)/T)+B\cos(10(2\pi t)/T)$ $\qquad+C\sin(3(2\pi t)/T)+D\cos(3(2\pi t)/T)$ $A,B,C,D\stackrel{{\scriptstyle ind}}{{\sim}}U(0,5)$ $y_{t}\sim\mathcal{N}(0,\exp(h_{t}))$ $j:=\left\lfloor\frac{t}{25}\right\rfloor+1$ $h_{t}=(-1)^{j}\cdot|z_{j}|$ $z_{j}\sim\begin{cases}N(5,0.5^{2})&\text{if $j$ is even}\\ N(0,0.5^{2})&\text{if $j$ is odd}\end{cases}$
Table 3: Summary of the eight data‐generating processes (DGPs). DGPs 1–3 simulate stochastic‐volatility models with one, two, or three regimes, and DGPs 4–6 simulate GARCH(1,1) models with one, two, or three regimes. In DGP 7, $h_{t}$ is a random sinusoid, and in DGP 8, $h_{t}$ is piecewise constant with random intercepts. For DGPs 2 and 4, the transition probability of remaining in the same state is 0.98, and transitioning to a different state is 0.02. For DGPs 3 and 6, the probability of staying in the same state is 0.98, while switching to any other state occurs with probability 0.01. Each DGP comprises 1,000 simulated sample paths, each consisting of 1,000 observations.